7. The standard deviation of the mean for the sampling distribution of random samples of size n-36 from a large (or infinite) population is 2. How large must the sample size be to decrease the standard deviation to 1.2?
Answer #1

Given that, the standard deviation of the mean for the sampling distribution of random samples of size n = 36 from a large population is 2. That is, 2

Therefore, the population standard deviation (\sigma)

\sigma_{\bar x } = \frac { \sigma}{\sqrt { n }}

=> \sigma= \sigma_{\bar x } *\sqrt { n } = 2* \sqrt { 36}= 12

We want to find, the sample size be to decrease the standard deviatiom to 1.2

=>n = (\frac { \sigma}{ \sigma_{\bar x } })^2= (\frac { 12}{1.2})^2 = 100

Required sample size is 100